Julio borrows $250 from Ricky. Ricky wants to make a 5% real return on his money, so they both agree on a 5% interest rate paid next year. Both don't anticipate the 5% inflation next year. In this case

a) Julio is better off.

b) Ricky will receive more than 5% of real rate of return a year from now.

c) Ricky is better off.

d) Julio will pay $15 a year from now on.

The CPI somewhat overstates changes in the cost of living because it does not allow for substitutions that consumers might make in response to price changes.

a) True

b) False

An inflation rate that is lower than expected benefits creditors.

a) True

b) False

Which of the following is not a desirable characteristic in an economy?

a) low unemployment

b) population growing slower than output

c) rapid increase in the general price level

d) growing per-capita output

If the consumption function is of the form [C = 80 + 0.4Y], the MPS equals

a) 0.4.

b) 0.6.

c) -0.6.

d) -0.4.
Answer #1

(1) (a)

Unexpected inflation reduces real value of loan. Since loan is liability for borrower, the borrower gains from inflation.

(2) True

Substitution bias causes CPI to overstate inflation.

(3) True

Lower inflation increases real value of loan. Since loan is asset for lender, the lender gains from lower inflation.

(4) (c)

Rapidly rising price level slows down economic growth.

(5) (b)

MPC = 0.4

MPS = 1 - MPC = 1 - 0.4 = 0.6